OpenStudy (anonymous):

Find an equation of the line that is tangent to the graph of f and parallel to the given line f(x)=3x^3, 9x-y+9=0 I need help doing this one

OpenStudy (perl):

y = 9x + 9 , this has slope 9 f ' (x) = 9x^2 solve f '(x) = 9 9x^2 = 9 x^2 = 1 x = 1 , -1